For the chemical reaction,

2HCl + Ca(OH)2 -> CaCl2 + 2H2O

What mass of calcium hydroxide in grams is needed to produce 3.71 mol of water?
Answer #1

2HCl +Ca(OH)2 - - > CaCl2 +2H2O

as here it is clear that

1 mole of Ca(OH)2 will produce 2 mole of H2O.

molecular weight of Ca(OH)2 = 40 +16+16+2 =74 g/mol

molecular weight of H2O = 2+16 =18g/mol

1 mole of Ca(OH)2 contains 74 g of Ca(OH)2.

As 2 mole of water is produced by 1 mole of Ca(OH)2.

or

2Mole of water is producef by 74g of Ca(OH)2.

then 3.71 mole of water is produced by 74*3.71/2 =137.87 g of Ca(OH)2.